I1) Consider a population with data values of 12 8 28 22 12 30 14. The population mean is B) 12 C) 18. D) 14.
Answer #1

11)

population Mean =\frac{\sum{Xi}}{N}

population Mean = \frac{12+8+28+22+12+30+14}{7}

population Mean =\frac{126}{7}=18

option C is correct

c. 18